A train moves 3.6km from rest to rest in 3 minutes. The greatest speed is 90km/h, and the acceleration and retardation are uniform. Find the distance traveled at full speed.
1
Expert's answer
2021-08-18T11:08:36-0400

Given,

Distance (d)= 3.6 km

"(d)= 3.6\times 1000 m = 3600m"

Time "(t) = 3min = 3\times 60 = 180s"

greatest speed "(v)=90 km/h = 90\times \frac{1000}{60\times 60} m/s"

"\Rightarrow (v)=\frac{15\times 10}{6}m/s = 25 m/s"

initial speed "(u) = 0", let the acceleration of the train be a, and time during the acceleration be "t_1" and time for the de-acceleration be "t_2"

Motion during acceleration,

"\Rightarrow v=u+at_1"

"\Rightarrow 25=at_1"

"\Rightarrow t_1 = \frac{25}{a}"

"d_1 = \frac{at_1^2}{2}=\frac{625}{2a}"

motion during de-acceleration,

"\Rightarrow v_f=v-at_2"

Now, substituting the values,

"\Rightarrow 0=25-at_2"

"\Rightarrow at_2 = 25"

"\Rightarrow t_2=\frac{25}{a}"

"v_f^2=v^2-2ad_2"

"\Rightarrow 0= 25^2-2ad_2"

"\Rightarrow d_2=\frac{625}{2a}"

Time, during which train traveled with constant velocity,

"d_3 = 25t_3"

given that,"t_1+t_2+t_3 = 180s"


"\Rightarrow t_3= 180-(t_1+t_2)"


so, "d_3=(180-(t_1+t_2))\times 25"


"\Rightarrow d_3 = (180-(\frac{25}{a}+\frac{25}{a}))\times 25"


"\Rightarrow d_3= (180- \frac{50}{a})\times 25"


Now, "d_1+d_2+d_3 = 3600m"


"\Rightarrow \frac{625}{2a}+\frac{625}{2a}+(180-\frac{50}{a})\times 25 = 3600m"


"\Rightarrow \frac{1250}{2a}+4500-\frac{1250}{a}=3600"


"\Rightarrow \frac{1250}{a}-\frac{1250}{2a}=4500-3600"


"\Rightarrow \frac{1250}{2a}=900"


"\Rightarrow a =\frac{1250}{1800}=0.7m/s^2"


now, "t_1=t_2 = \frac{50}{0.7}" sec


"\Rightarrow t_1=t_2 = 71.42 sec"


now "d_3 = (180-71.42-71.42)\times 25"

"\Rightarrow d_3 = 929m"

Hence, the distance traveled with full speed is "929m"